OpenStudy (anonymous):

how many edges does a polyhedron have , that has 12 faces

OpenStudy (anonymous):

help , please

OpenStudy (anonymous):

Alrighty, The formula for finding the number of edges is V - E + F = 2.

OpenStudy (anonymous):

V = 2 + 12 - 6 = 8.